# V2XPnP: Vehicle-to-Everything Spatio-Temporal Fusion for Multi-Agent Perception and Prediction

### v1

@@ 5.2 Benchmark Methods
End-to-end methods. We focus on the end-to-end models in this paper and implement an end-to-end model for cooperative temporal perception and prediction. This model uses the same LiDAR backbone and decoding heads with V2XPnP but utilizes a simple temporal fusion module FaF [29] - alternating 2D and 3D convolutions- as the No fusion-FaF baseline. Then, the Early fusion configurations and two state-of-the-art single-frame intermediate fusion models, V2VNet [35] and F-Cooper [6] to integrate the end-to-end model as the early fusion and intermediate fusion baselines. These models marked with are reimplemented in our framework to ensure consistency in their LiDAR backbone, temporal fusion network, and decoding heads.
Decoupled methods. Transmitting final detection results rather than spatio-temporal features renders late fusion incompatible with end-to-end models. Therefore, we benchmark Late Fusion with a decoupled pipeline, where objects in each historical frame are detected using a single-frame perception module, and results are fused via non-maximum suppression [42]. Assuming an ideal tracker to generate object trajectories from perception results and interpolate missing points, we implement an attention-based predictor for trajectory-level prediction tasks, following the prediction model mainstream [33]. To further assess end-to-end performance, we also evaluate a decoupled No-Fusion model.
@@ 5.3 Results
Tab. 1 presents the benchmark results across four V2X collaboration modes. Since prediction performance is inherently tied to detection quality, the difficulty posed by different detected objects can significantly impact prediction accuracy. As a result, the EPA metric serves as the most appropriate indicator for assessing the overall performance across models. More details and baseline results are provided in the supplementary material.
Transmit. In both end-to-end and decoupled PnP frameworks, the cooperation perception and prediction performance are consistently better than non-fusion models in all collaboration modes, especially in the primary EPA metric, which demonstrates the benefits of cooperation in temporal perception and prediction. Moreover, intermediate fusion models (e.g., F-Cooper and V2XPnP) generally outperform other fusion strategies, while early fusion consistently surpasses late fusion. However, V2VNet has a worse performance, likely due to the absence of a map and the loss of temporal features during explicit feature ROI matching. Our V2XPnP model archives the best performance, with an average improvement of 15% and 24% over F-Cooper in AP and EPA, respectively, because the unified Transformer model better captures the complex relationship among the spatio-temporal information shared by V2X.
@@
When to Transmit. Tab. 2 shows the performance of our model under two communication strategies. (1) The perception and prediction performance of one-step communication is improved compared to multi-step communication without delay by 5.0% AP and 12% in EPA. This improvement arises because the ego agent directly fuses lossless raw data, avoiding error accumulation from lossy intermediate information transformed across temporal dimensions. (2) The one-step strategy compresses spatio-temporal feature transmission (under a 32$\times$ compression rate) from 5$\times$0.269 Mb to 0.269 Mb compared to multi-step communication, while mitigating information loss when agents move out of communication range in historical frames. At a typical C-V2X data transmission rate [2], the transmission delay of one-step communication is approximately 10$\sim$20 ms. (3) Under the noise settings from [41], we provide the results with 100 ms time delay and positional/head Gaussian noise (0.2m, 0.2°). Due to error accumulation during multiple transmissions, the performance of the multi-step communication strategy decreases by 14%, while the one-step strategy’s performance only decreases by 3.7%, demonstrating greater robustness.
Table 2: Comparison of one-step and multi-step communication; Bold values indicate results under delay and noisy settings.
@@
68.5 -4.3% 1.37 2.38 31.6 46.4 -3.7%
End-to-end vs. Decoupled Frameworks. The end-to-end model consistently outperforms the decoupled framework. In no-fusion situations, the end-to-end FaF model outperforms the decoupled model in detection by leveraging temporal cues. Furthermore, FaF achieves performance comparable to late fusion with V2X spatial aggregation due to integrating temporal features. The intermediate fusion of spatio-temporal features in V2XPnP aligns well with the end-to-end architecture, showcasing its superior performance.
Infrastructure vs. Vehicle Centric. As shown in Tab. 1, models under VC and IC modes outperform those in V2V and I2I modes, because VC and IC can aggregate information from up to four agents rather than only two, resulting in enhanced environmental understanding. Additionally, stationary infrastructure-based agents in IC and I2I modes offer higher prediction accuracy by providing elevated sensing perspectives and less noisy sensor data.
Qualitative Results. Fig. 5 visualizes the outcomes of cooperative perception and prediction across different fusion models. The No Fusion model is constrained by its limited field of view and thereby shows subpar performance. FaF*, leveraging temporal information within an end-to-end pipeline, performs better under occlusion. Late and early fusion models significantly benefit from multi-agent data integration, though late fusion remains impacted by error propagation from perception, such as heading errors misleading trajectory direction. Notably, the end-to-end intermediate fusion model, particularly V2XPnP, achieves superior performance in both detection and prediction tasks.
@@ 6 Conclusions
We propose V2XPnP, a novel V2X spatio-temporal fusion framework for cooperative temporal perception and prediction. The core of this framework is a unified Transformer-based model for spatio-temporal fusion and map fusion. Furthermore, we examine various fusion strategies concerning what, when to transmit, and how to fuse, offering comprehensive benchmarks. Additionally, we introduce the V2X Sequential Dataset, which supports all V2X collaboration modes. Extensive experimental evaluations demonstrate the superior performance of the proposed framework, establishing its efficacy in advancing V2X-enabled cooperative temporal tasks. Future work will focus on developing more effective fusion models and adaptive communication strategies that dynamically optimize bandwidth utilization.
@@ A.1 Benchmark Model Details
Intermediate Fusion Methods. We implement two state-of-the-art single-frame intermediate fusion methods (i.e., V2VNet [35] and F-Cooper [6]) and extend them to spatio-temporal fusion using a baseline temporal fusion model that alternates between 2D and 3D convolutions, denoted as V2VNet and F-Cooper. The model setting and configurations in multi-agent fusion follow the original implementation in [35] and [6]. Specifically, V2VNet employs direct max-pooling, while F-Cooper utilizes spatial-aware message-passing mechanisms to aggregate features across multiple agents. The feature map dimensions in our experiments are $176\times 100$ with $256$ channels.
@@ A.4 Training Strategy
Stage 2: Multi-Agent Spatio-Temporal Learning. Based on the pre-trained single-agent model, the multi-agent fusion module is introduced and jointly trained with the entire model. At this stage, the primary focus is to balance the two tasks, ensuring that neither perception nor prediction dominates the training process. To achieve this, we employ a dynamic loss-weighting strategy that gradually increases the weight assigned to the prediction loss. This approach ensures balanced optimization, avoiding performance trade-offs between tasks and improving overall effectiveness across both perception and prediction objectives.
Training Details. The model is trained using the Adam optimizer with an initial learning rate of $2\times 10^{-3}$ and a weight decay of $1\times 10^{-4}$. We employ $4$ training stages detailed before, and each training stage consists of $20$ epochs with a batch size of $2$, and early stopping is employed to prevent overfitting. We carefully tune the hyperparameters to ensure the stability and efficiency of the training process.
@@
B Cooperative Temporal Perception Task
In addition to the end-to-end perception and prediction task, the sequential nature of our V2XPnP-Sequential dataset facilitates other temporal tasks, including temporal perception and traditional prediction tasks. In this section, we introduce the cooperative temporal perception task and present benchmark results on the V2XPnP-Sequential dataset. Details on the traditional prediction task are provided in Sec. C.
B.1 Problem Formulation
B.2 Benchmark Methods
B.3 Benchmark Results
C Traditional Cooperative Prediction Task
C.1 Problem Formulation
@@ C.3 Benchmark Results
D V2XPnP Sequential Dataset Details
D.1 Dataset Visualization
D.2 Data Acquisition
@@
D.3 Data Annotation and Processing
Data Annotation. The 3D bounding boxes in our V2XPnP-Sequential dataset are annotated using an open-source labeling tool, SUSTechPOINTS [23], by expert annotators. The first step is annotating the bounding boxes in the point clouds from the two CAVs and infrastructure units. Then, these bounding boxes, annotated in different agents’ coordinate frames, are processed through a V2X sequential pipeline to assign consistent object IDs across agents and temporal frames. To ensure annotation quality, each object is subjected to eight rounds of review and revision. In total, ten object categories are included in our dataset: car, pedestrian, scooter, motorcycle, bicycle, truck, van, concrete truck, bus, and road barrier. Each object annotation includes the center of the bounding box ($x,y,z$), sizes ($width,length,height$), and orientation ($roll,yaw,pitch$) in the global coordinates.
@@ D.4 Dataset Analysis
D.5 Dataset Privacy Protection

### v2

@@ 5.2 Benchmark Methods
End-to-end methods. Most of the single-frame perception models cannot support the prediction task, thus, we first implement a baseline end-to-end model with the same LiDAR backbone and decoding heads as V2XPnP but utilizing the temporal fusion module FaF [33] - alternating 2D and 3D convolutions - as the No fusion-FaF baseline, which can extend the non-temporal model to support temporal tasks. Then, the early fusion configurations and several state-of-the-art intermediate fusion models are integrated as: Early Fusion, FFNet [51], CoBEVFlow [42], V2X-ViT [45]). More benchmark results of DiscoNet [27], F-Cooper [6], and V2VNet [40] are provided in the supplementary. These models marked with are reimplemented in our framework with the same LiDAR backbone and decoding heads.
Decoupled methods. Transmitting final detection results renders late fusion incompatible with end-to-end models. Thus, we benchmark Late Fusion with a decoupled pipeline, where objects in each historical frame are detected using a single-frame perception module, and results are fused via non-maximum suppression. Assuming an ideal tracker to generate object trajectories from perception results and interpolate missing points, we implement an attention-based predictor for trajectory-level prediction tasks, following the prediction mainstream [37]. To further assess end-to-end performance, we also evaluate a decoupled No-Fusion model.
@@ 5.3 Results
Tab. 1 presents the benchmark results across four V2X collaboration modes. Since prediction performance inherently depends on detection quality, the difficulty posed by different detected objects can significantly impact prediction accuracy. Thus, the EPA metric serves as the most appropriate indicator for assessing the overall performance. More details and baseline results are provided in the supplementary.
What to Transmit. In both end-to-end and decoupled PnP frameworks, the cooperation perception and prediction performance are consistently better than non-fusion models in all collaboration modes, especially in the primary EPA metric, which demonstrates the benefits of cooperation in temporal perception and prediction. Moreover, intermediate fusion models (e.g., V2X-ViT, FFNet, and V2XPnP) generally outperform other fusion strategies, while early fusion consistently surpasses late fusion. Our proposed V2XPnP model achieves the best performance, outperforming other competitive cooperative methods.
@@
Table 2: Comparison of one-step and multi-step communication
@@
When to Transmit. Tab. 2 shows the performance of our model under two communication strategies. We find that the perception and prediction performance of one-step communication is improved compared to multi-step communication by 5.0% AP and 12% in EPA. This improvement arises because each agent first directly fuses their lossless raw temporal data before sharing, avoiding error accumulation from lossy intermediate information transformed in multi-step communication across temporal dimensions. Moreover, the one-step strategy compresses spatio-temporal feature transmission (under a 32$\times$ compression rate) from 5$\times$0.269 Mb to 0.269 Mb compared to multi-step communication, while mitigating information loss when agents move out of communication range in historical frames. At a typical C-V2X data transmission rate [4], the transmission delay of one-step communication is approximately $10\sim 20$ ms.
End-to-end vs. Decoupled Frameworks. The end-to-end model consistently outperforms the decoupled framework. In no-fusion situations, the end-to-end model FaF outperforms the decoupled model in detection by leveraging temporal cues. Furthermore, FaF achieves performance comparable to late fusion with V2X spatial aggregation due to integrating temporal features. The intermediate fusion of spatio-temporal features in V2XPnP aligns well with the end-to-end architecture, showcasing its superior performance.
Infrastructure vs. Vehicle Centric. As shown in Tab. 1, models under VC and IC modes outperform those in V2V and I2I modes, because VC and IC can aggregate information from up to four agents rather than only two, resulting in enhanced environmental understanding. Notably, the evaluation protocol is consistent across all models within each mode. Additionally, stationary infrastructure-based agents in IC and I2I modes offer higher prediction accuracy by providing elevated sensing perspectives and less noisy data.
Transmission Data Size and Robustness Test. Results in Fig. 6 indicate that V2XPnP achieves a good balance between communication efficiency and accuracy, maintaining superior performance compared to full-size V2X-ViT* even at a ×$\times 128$ compression rate. Following the setting in [45], we provide the results with 100-500 ms time delay and positional/head Gaussian noise from (0.2m, 0.2°) to (1m, 1°). Both V2XPnP and V2X-ViT* maintain robust performance due to their spatial attention fusion, and V2XPnP performs better due to designed temporal attention.
Qualitative Results. Fig. 5 visualizes the outcomes of cooperative perception and prediction across different fusion models. The No Fusion model is constrained by its limited field of view. The FaF model, leveraging temporal information within an end-to-end pipeline, performs better under occlusion. Late and early fusion models significantly benefit from multi-agent data integration, though late fusion remains impacted by error propagation, such as detection heading errors misleading trajectory direction. Notably, the end-to-end intermediate fusion model, particularly V2XPnP, performs better in both detection and prediction tasks.
Figure 6: Transmission data size and communication noise test results. V2XPnP shows better performance with varying data compression rates and robustness under communication noise.
@@ 6 Conclusions
We propose V2XPnP, a novel V2X spatio-temporal fusion framework for cooperative temporal perception and prediction. The core of this framework is a unified Transformer-based model for spatio-temporal fusion and map fusion. Furthermore, we examine various fusion strategies concerning what, when to transmit, and how to fuse, offering comprehensive benchmarks. Additionally, we introduce the V2X Sequential Dataset, which supports all V2X collaboration modes. Extensive experiments demonstrate the superior performance of the proposed framework, establishing its efficacy in advancing cooperative temporal tasks.
@@ A.1 Benchmark Model Details
Intermediate Fusion Methods. We implement several state-of-the-art single-frame intermediate fusion methods, including V2VNet [40], F-Cooper [6], DiscoNet [27],CoBEVFlow [42], FFNet [51], V2X-ViT [45], and our proposed V2XPnP model, integrating them with our end-to-end model to replace the spatio-temporal fusion module. The model settings and configurations for the fusion module adhere to the original implementations.
@@ A.4 Training Strategy
Table S1: Additional benchmark results of cooperative perception and prediction models on V2XPnP Sequential (V2XPnP-Seq) Dataset
V2XPnP-Seq-VC (with V+2I at most) V2VNet* [40] 48.6 2.10 3.75 42.3 25.3
F-Cooper* [6] 66.0 1.35 2.56 36.1 38.7
DiscoNet* [27] 66.8 1.41 2.62 34.4 42.8
V2XPnP (Ours) 71.6 1.35 2.36 31.7 48.2
V2XPnP-Seq-IC (with 2V+I at most) V2VNet* [40] 33.6 1.95 3.53 44.2 16.3
F-Cooper* [6] 60.2 1.21 2.32 36.3 36.3
DiscoNet* [27] 65.4 1.14 2.18 36.1 40.7
V2XPnP (Ours) 71.0 1.18 2.16 34.0 46.0
V2XPnP-Seq-V2V V2VNet* [40] 43.1 3.10 5.55 46.8 19.4
F-Cooper* [6] 60.2 1.69 3.22 41.1 34.4
DiscoNet* [27] 61.2 1.66 3.13 41.2 33.1
V2XPnP (Ours) 70.5 1.78 3.28 39.9 40.6
V2XPnP-Seq-I2I V2VNet* [40] 41.1 1.83 3.34 40.4 23.2
F-Cooper* [6] 58.6 1.34 2.58 40.0 33.6
DiscoNet* [27] 63.5 1.15 2.19 37.5 38.4
V2XPnP (Ours) 69.2 1.26 2.31 36.5 42.8
Stage 2: Multi-Agent Spatiotemporal Learning. Based on the pre-trained single-agent model, the multi-agent fusion module is introduced and jointly trained with the entire model. At this stage, the primary focus is to balance the two tasks, ensuring that neither perception nor prediction dominates the training process. To achieve this, we employ a dynamic loss-weighting strategy that gradually increases the weight assigned to the prediction loss. This approach ensures balanced optimization, avoiding performance trade-offs between tasks and improving overall effectiveness across both perception and prediction objectives.
Training Details. The model is trained using the Adam optimizer [22] with an initial learning rate of $2\times 10^{-3}$ and a weight decay of $1\times 10^{-4}$ with early stopping on NVIDIA L40S GPUs. We employ $4$ training stages, as detailed before, and each training stage consists of $30$ epochs with a batch size of $2$. Early stopping is employed to prevent overfitting. We carefully tune the hyperparameters to ensure the stability and efficiency of the training process.
@@
B Additional Benchmark Results
In this paper, we benchmark different spatiotemporal strategies with 11 fusion models in total:
  • No Fusion: No Fusion, No Fusion-FaF
  • Early Fusion: Early Fusion
  • Late Fusion: Late Fusion
  • Intermediate Fusion: V2VNet [40], F-Cooper [6], DiscoNet [27], CoBEVFlow [42], FFNet [51], V2X-ViT [45], and our proposed V2XPnP.
We present additional benchmark results for V2VNet [40], F-Cooper [6], and DiscoNet [27] across all collaboration modes, as shown in Tab. S1. Our proposed V2XPnP consistently outperforms these SOAT baselines in terms of EPA and AP across all collaboration modes. Notably, V2VNet* exhibits lower performance, likely due to the absence of a map and the loss of temporal features during explicit feature ROI matching.
C Cooperative Temporal Perception Task
In addition to the end-to-end perception and prediction task, the sequential nature of our V2XPnP-Sequential dataset facilitates other temporal tasks, including temporal perception and traditional prediction tasks. In this section, we introduce the cooperative temporal perception task and present benchmark results on the V2XPnP-Sequential dataset. Details on the traditional prediction task are provided in Sec. D.
C.1 Problem Formulation
C.2 Benchmark Methods
C.3 Benchmark Results
@@
D Traditional Cooperative Prediction Task
D.1 Problem Formulation
@@ D.3 Benchmark Results
E V2XPnP Sequential Dataset Details
E.1 Dataset Visualization
E.2 Data Acquisition
@@
E.3 Data Annotation and Processing
Data Annotation. The 3D bounding boxes in our V2XPnP-Sequential dataset are annotated using an open-source labeling tool, SUSTechPOINTS [25], by expert annotators. The first step is annotating the bounding boxes in the point clouds from the two CAVs and infrastructure units. Then, these bounding boxes, annotated in different agents’ coordinate frames, are processed through a V2X sequential pipeline to assign consistent object IDs across agents and temporal frames. To ensure annotation quality, each object is subjected to eight rounds of review and revision. In total, ten object categories are included in our dataset: car, pedestrian, scooter, motorcycle, bicycle, truck, van, concrete truck, bus, and road barrier. Each object annotation includes the center of the bounding box ($x,y,z$), sizes ($width,length,height$), and orientation ($roll,yaw,pitch$) in the global coordinates. Notably, we follow a general object definition in annotation, encompassing stationary objects such as parked vehicles and barriers, which are annotated similarly to movable objects but explicitly labeled as static. This aligns with public datasets like nuScenes [5], where static objects are tracked while maintaining consistent IDs.
@@ E.4 Dataset Analysis
E.5 Dataset Privacy Protection